COUPLED CLUSTER CALCULATIONS FOR MOLECULES IN TRIPLET GROUND STATES

\underline{P. BOTSCHWINA}, Institut für Physikalische Chemie, Universität Göttingen, Tammannstraße 6, D-37077 Göttingen, Germany.

On the basis of large-scale open-shell coupled cluster calculations accurate equilibrium structures are established for a number of molecules in triplet ground states. Among these are species of type HC_{2n}N (n = 1-3), C_{2n}S (n = 1-3), SiC_{2n+1} (n = 1-3) and SiC_{2n−1}S (n = 1-4). The equilibrium structure of HC_{2}N is definitely non-linear with a barrier height to linearity close to 300 cm^{−1}. On the other hand, triplet HC4N has a linear equilibrium structure and this probably also holds for the larger members of the series. Various spectroscopic properties calculated for the above mentioned molecules are discussed.